QUESTION 5

A bus company advertised a mean time of 150 minutes for a

trip between two cities. A consumer group had reason to

believe that the mean time was more than 150 minutes. A

sample of 40 trips showed a mean ˉx=153 minutes and a

standard deviation s=7.5 minutes. At the 0 .05 level of

significance, test the consumer group’s belief.
Answer #1

\\Given,\\ Level~of~significance(\alpha)=0.05\\\\ sample~Size(n)=40\\\\ Standard~Deviation(s)=7.5\\\\ Sample~Mean(\bar{x})=153\\\\ .....................................................\\ Null~Hypothesis~H_0:\mu=150\\\\ Alternative~Hypothesis~H_1:\mu>150\\\\ ........................................................\\ Test~Statistic(t)=\frac{\bar{x}-\mu}{s/\sqrt{n}}=\frac{153-150}{7.5/\sqrt{40}}=2.53\\\\ .........................................................\\ Degree~of~freedom=n-1=40-1=39\\\\ P-value=P(t>2.53)\\\\ ~~~~~~=0.0078\\\\ (use~excel~function:=TDIST(2.53, 39,1))\\\\ ............................................................\\\\ as~p-value~is~less~than significance~level(\alpha=0.05)~Reject~Null~Hypothesis.\\\\ ............................................................\\ Conclusion:Hence, ~there~is~Enough~evidence~to~support~the~claim~that~the~mean~time~was~more~than~150~minutes.